OpenStudy (anonymous):

What is the y-intercept for the graph of 3x – 5y = 6?

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

ganeshie8

OpenStudy (anonymous):

Heyy this is the spanish section :)

OpenStudy (anonymous):

But ... 3 x 0 - 5y = 6 -5y = 6 y = -6/5

OpenStudy (anonymous):

yeah, why would you type this in the Spanish section?